### v1

# From Wavefunctional Entanglement to Entangled Wavefunctional Degrees of Freedom

Abstract
The question of whether entanglement between photons is equivalent to entanglement between their characteristic field modes—specifically, the single-particle wavefunctions that are composed and superposed to describe identical particles—is a key, open problem concerning multi-partite optical degrees of freedom, and has profound implications for topics ranging from quantum foundations to quantum computation. In this Letter, I offer a fresh, deeper, physical insight into this subtle, albeit enduring, issue by describing a situation in which entangling interactions between optical modes—namely, the wavefunctions—can be distilled into genuine entanglement between the physical properties of the photons—which are the wavefunctional degrees of freedom. This theoretical observation also highlights the salience of the measurement context—especially, of clearly disambiguating between the choice of the quantum subsystem and the decision to measure an observable along a particular axis of measurement—while quantifying and transforming quantum optical entanglement. This understanding might be applied to formulate a new class of protocols for distilling quantum magic from contextually and nonlocally entangled photons within inseparable field modes.
Introduction.—Quantum entanglement and quantum nonlocality [1, 2, 3, 4, 5, 6, 7, 8] are fundamental features of the quantum mechanical description of many-particle systems, and play defining roles in foundations of quantum physics [9, 10, 11, 12]; quantum information science and engineering [13, 14, 15, 16]; atomic [17, 18, 19, 20], molecular [21], optical [22, 23, 24, 25, 26, 27], condensed matter [28, 29], and high-energy physics [30]; as well as in cosmology [31, 32, 33, 34]. While the earliest intimations of the notion of quantum nonlocality are traceable to the 1927 Solvay Conference [*[See, e.g., ][forahistoryofthoughtonentanglement, nonlocality, andincompleteness.]Bacciagaluppi_Crull_2024, 36, 37], in 1935, Einstein, Podolsky, and Rosen (E.P.R.) [1] formally introduced the concept of quantum entanglement in the course of devising a Gedankenexperiment to clearly show a conflict between: the quantum mechanical postulate that the wavefunction provides a complete description of the physical reality of a quantum system; and the absence of simultaneous elements of physical reality corresponding to values of mutually complementary physical quantities that are described by non-commuting operators. Specifically, E.P.R. argued that either or is true, but not both. While, in the standard quantum theory, is true and is false, E.P.R. were able to demonstrate that the situation in which the coordinates, as well as the momenta of two particles are entangled, paradoxically suggests that is false, but is true.
A more careful reading of their paper reveals that E.P.R. devised entanglement as a mechanism to predict the value of a physical quantity of a quantum system, say, system I with “certainty” [1]—that is, with probability equal to unity, like in a classical, realist theory—without in any way “disturbing” [1]—that is, altering the physical state of—the system. The actual, direct measurement were to be carried out on the system’s entangled, spatially separated counterpart, say, system II. Crucially, E.P.R. implicitly assumed that the principle of classical locality should ensure that such an indirect measurement on system II would not “disturb” system I [2, 38].
Background.—Guided by Bohm’s reformulation of the E.P.R. paradox [39] in terms of discrete-valued, dichotomous observables, namely, spin angular momenta, Bell correctly realized that the above-described conflict is actually one between the irrevocable indeterminism of quantum mechanics and E.P.R.’s presupposed, cherished notion of local, classical realism [40, *bell2004speakablepaper2]. Simply put, situations, such as the E.P.R.B. (E.P.R.-Bohm) entanglement in which there are perfect correlations—or anti-correlations—between the outcomes of measurements performed upon two causally disconnected, space-like separated particles suggest that there are two opposing descriptions of quantum reality, namely, either: physical reality is non-local, measurement outcomes are inescapably probabilistic, and the wavefunction description of physical reality is complete; or physical reality is local, measurement outcomes are predetermined by hidden elements of reality, and therefore the wavefunction description of physical reality is incomplete, which is the crux of the E.P.R. argument [42, *werner2014comment, *wiseman2014two].
Notably, Bell resolved the E.P.R. paradox by formulating empirically verifiable inequalities [40, *bell2004speakablepaper2] that quantified the difference between and above. Bell’s inequalities bound from above the strengths of correlations between the outcomes of measurements on two quantum systems in situations in which each of such systems has its own, individual properties, as is suggested by local realist theories that are formulated in the spirit of the E.P.R. viewpoint, like above.
A series of exceedingly elegant and increasingly loophole-free experiments, which were carried out between 1972 and the present time—see, e.g., Refs. [22, 23, 24, 25, 26, 45, 46, 47, 48, 49, 50], and references therein—conclusively demonstrated an explicit violation of refined versions of Bell’s inequalities [51, 52, 53], such as the Bell-C.H.S.H. inequalities [51]. These experiments are enormously significant, as they decisively rule out any conceivable, deterministic, local realist, hidden-variables theory [54, *bell2004speakablepaper1] that attribute the statistical nature of measurement outcomes to the underlying statistical distributions of hidden, supplementary elements of physical reality, analogous to classical probability distributions of classical random variables.
Strikingly, Greenberger, Horne, and Zeilinger (G.H.Z.) extended Bell’s results by showing that—under special circumstances—the definite, certain predictions about measurement outcomes by standard quantum theory directly conflict the corresponding predictions by local realist theories, as opposed to a mere statistical contradiction involving inequalities based on correlation functions [56, 57]. Such situations are realized by probing specific kinds of multi-particle entangled states, known as the G.H.Z. states [58, 59, 27]. -particle G.H.Z. states, where , are commonly used in optical quantum computing schemes and protocols [60, 61, 62, 63, 64, 65, 66, 67, 68, 69] that, fundamentally, derive their quantum advantage from the intrinsic quantum nonlocality of such highly entangled states [70, 60]. More broadly, scalable, quantum computational advantage has been shown to originate from quantum contextuality, which is a generalized notion of quantum nonlocality [71]. Moreover, the quantum mechanical predictions for these multi-particle entangled states are completely independent of the relative arrangements of the measuring apparatuses in space and time.
Motivation for this work.—The quantum mechanical systems of choice in a majority of the above-described experiments were optical, which raised an accompanying, profound question: (Qu1) Is entanglement between modes of light equivalent to entanglement between individual particles of light? (see, e.g., Refs. [72, 73, 74] for a detailed discussion on the history of, and related work on this key problem.) Resolution of this quantum foundational issue also has significant implications for quantum information processing [75] and quantum-enhanced sensing [76] with highly entangled states of light.
Usually, in quantum optics, a bosonic mode is modeled as a single-particle wavefunction, either in real space, or in some other space, such as momentum space [77]. This description, however, does not lend itself to direct and unequivocal correspondence with observables, as opposed to a description based on photons, which are fundamental particles, and, therefore, have unambiguously real properties. For example, a beam of light can be described by multiple equivalent—and all mathematically legitimate—bases of modes, even though one, or a preferred set of such mode bases might be particularly convenient for describing and calculating the properties of the light. This feature leads to the question of the relative nature of entanglement between modes, namely: (Qu2) If one changes to a new mode basis, does the inter-modal entanglement still exist? In contrast, inter-particle entanglement is invariant under basis transformations.
As is well known, protocols for deterministically preparing—that is, without requiring any quantum measurements at any step—entangled photonic states—such as the state, , where and represent two spatial modes of photon 1 (quantum subsystem 1), and are two spatial modes of photon 2 (quantum subsystem 2), and and are suitable normalizing, expansion constants—require either direct or effective interactions between photons, both of which are unusually difficult to realize in practice. On the one hand, coherent, entangled superpositions of light field modes are readily prepared by amplitude and beam splitting, as well as by nonlinear optical processes [78]; recently, for example, an expanded set of such superpositions has been shown to be deterministically realizable by higher-dimensional, noiseless, quantum holonomic approaches [74]. On the other hand, the pioneering experiments of Zeilinger and his colleagues have shown that quantum measurements play an important role in the creation of G.H.Z. states; for example, the detection of an ancillary, trigger photon heralds the formation of entanglement between photons.
Consequently, one may pose the question, which seems both fundamentally, as well as practically important: (Qu3) Is it possible to transfer the inter-modal entanglement—which is relatively simpler to achieve—to inter-particle entanglement—which is particularly challenging to realize—as deterministically as possible? The purpose of this Letter is to show that, indeed, wavefunctional 111Throughout this Letter, the term wavefunctional is used as the adjectival form of wavefunction, and is unrelated to a wave functional that plays the role of a wavefunction in a quantized theory of wave fields [93]. entanglement—that is, entanglement between mathematically inseparable quantum wavefunctions or optical field modes—can be transformed into direct entangling interactions between optical photons that are described by such modes. An additional step that entails making a measurement on an ancillary photon appears to be irreplaceable. Figure 1 summarizes the problem that this Letter addresses, whereas Fig. 2 furnishes a proposed solution.
Disambiguating between choices of subsystems and of observables.—Before tackling question (Qu3), let us consider experiments of the kind that were pioneered by Aspect et al. [23, 25, 24] for demonstrating violations of Bell’s inequalities, so as to illustrate the role played by the relative context of the quantum measuring arrangements—specifically, that of local basis transformations on individual subsystems—in determining the specific definitions—or identities—of subsystems and structures of observables.
In such experiments, there is a source Q, which emits a pair of entangled photons, whose angles of linear polarization are nonlocally correlated. The two photons propagate in opposite directions and each individually encounters an analyzer of linear polarization—for instance, a linearly polarizing, beam-splitting cube, made of a birefringent crystal—that perfectly transmits the parallel—along the angle of orientation of the analyzer, such as and —as well as the perpendicular, linearly polarized components along two distinct directions. Each of such optical paths ends with a single-photon detector, thereby correlating the position of the photon with its angle of linear polarization, just like a Stern-Gerlach magnet oriented in space correlates the position of the electron—upon a detector screen—with the component of the spin angular momentum along the direction of the magnetic field. The locations of measurements of these angle of linear polarization—which are the entangled physical quantities—of the photons, A and B are assumed to be separated by a space-like interval.
Suppose that Q emits the pure, nonfactorable, E.P.R. state:
where A and B represent the two quantum subsystems—which are the two distinguishable photons—and H and V represent the horizontal and vertical orientations of the linear polarization, respectively. As is described in the accompanying end matter (also see Fig. 3), the above two-particle entangled state can be unitarily transformed to the one below, by performing individual basis transformations upon each of the two subsystems:
where the single-particle states, and are obtained by unitarily transforming and , respectively, for a given orientation angle. Performing distinct, local, unitary transformations—albeit of the same kind, namely, rotations by different angles—on the two causally disconnected subsystems is valid, since can be used to recover all the well-known expressions for the single- and two-particle, detection probabilities; the coefficients of correlation of linear polarization; as well as the Bell-C.H.S.H. inequalities. The entropies of entanglement of and are identical, and constant, regardless of the relative orientation, .
Remarkably, a similar situation can be orchestrated for entanglement between modes that describe neutral, bosonic atoms; the entangled property is the linear momentum of such atoms [77]. Four output modes are interfered two-by-two in two distinct spatial locations—with individually controllable phases, and —so as to demonstrate Bell nonlocality. The entropies of entanglement are also independent of the relative output phase, , which plays the role of above. A key assumption of this work is optical modes can be treated like any other bosonic mode, e.g., describing cold atoms.
Intriguingly, for modes, which are in photon occupation number entangled relationships with each other, we encounter this issue: The decision to measure the entangled property along, as it were, a particular axis of measurement—or, more precisely, a specifically chosen basis—and the identities—or the definitions—of the individual quantum subsystems, upon which such quantum measurements will be made, are inextricably linked, even though the Bell’s inequalities have been properly reframed for dealing with situations involving continuous, external degrees of freedom [80, 81, 82, 83, 84, 77]. For example, basis transformations concomitantly alter the occupation number—the entangled property—and the definition of the modal creation operator—the identity of the subsystem. For the above-described case of atoms with interfering modes, an additional element of subtlety is introduced by linear momenta—unlike spin angular momenta—along distinct axes commuting with each other. I invite the specialist reader to peruse the end matter for a deeper dive into this issue.
With the above discussion in mind, we are now ready to address (Qu3). The principal findings are summarized in the following principle:
The Equivalence between Wavefunctional and Inter-particle Entanglement Principle. For a quantum optical system having spatial modes, and operating with input, unentangled photons, the wavefunctional entanglement—that is, the entanglement between the single-particle quantum wavefunctions, or modes—can be transformed into entangling interactions between the photons that are described by such modes.
A physicist’s intuitive proof.—While I have given a somewhat more formal proof in the Supplemental Material [85], here, I will outline a simple argument that provides a deeper physical intuition into attacking this formidable problem. Figure 2 schematically summarizes this physical argument. Inspired by Ref. [77], I will model these spatial, bosonic modes as bound states, as opposed to, for example, propagating Gaussian wavepackets. I am specifically assuming that such modes can be mapped onto the bound energy eigenmodes of a simple harmonic oscillator, and of its anharmonically perturbed analogues; a quantum mechanical particle can always, meaningfully, said to be placed in a bound state.
Consider the following input state that describes two modes, which are in a photon occupation number entangled relationship with each other:
where and represent the two optical modes that are mapped onto the two oscillator energy eigenmodes, and and label the two photons that are supposed to be distinguishable. The symbol, indicates that the subsystems on the two sides of this symbol are different; for example, the entangled subsystems on the left (right) hand side above are modes (particles). Assuming identical photons would have implied: (up to normalization constants)—a phenomenon known as entanglement due to symmetrization [86, 72]. As this case is trivial and not useful for applications, I am considering distinguishable photons occupying distinguishable modes, with additional photonic degrees of freedom—for instance, angular momenta—ensuring that the overall state is symmetrized. To be as consistent as possible, in this simple Gedankenexperiment, with how optical fields are quantized in the second quantization approach, let us assume that the single-photon states map only onto the single-quantum oscillator eigenmodes. The zero- and two-photon states are assumed to map onto distinct oscillator modes, which, presumably, can be adiabatically eliminated from this problem, because of their distinct eigenenergies. Consequently, as depicted in Fig. 2(b), let us focus on the middle term, which corresponds to two photons occupying two distinct modes.
Let us now deploy a quantum mechanical random process that we will utilize to create macroscopic quantum superpositions. To accomplish this, we can introduce an “ancillary” quantum mechanical system, namely, a photon (photon 3 in Fig. 2) propagating through the Young’s double-slit experimental setup [see Fig. 3(a)]. As is well known, the single photon—as it transits through this apparatus—interferes with itself, and, consequently, is described by the superposition: of having gone through the top and bottom slits. Moreover, assume the screen to be photo-conducting, and without a lower half—so as to further emphasize that, unlike in the classical case, the position on the screen, e.g., the upper half, where the photon is detected is uncorrelated with the slit of entry, e.g., the top slit. The screen performs a measurement of the position of the photon. Now—assuming ideal detection and quantum efficiencies—the overall state of the detector and amplifier system, which has access only to photons impinging on the upper half, can be written as the superposition: of having registered and not registered the single photon as a single photo-electron. Assuming that the regions of photo-detection and collection of photo-electrons are sufficiently macroscopic, the probability amplitude, akin to the expression for the transition probability amplitude of a macroscopically delocalized detector atom for this diffraction experiment [87]. describes the spectral representation of the electromagnetic field propagating from the source to the detector.
The probabilistically ejected photo-electron is amplified, thereby producing a photo-current that drives an anharmonic potential controller, which controls the anharmonicity parameter, ; e.g., no photo-current corresponds to the initially harmonic potential. Consequently, the mode wavefunctions are now described as macroscopic superpositions of the energy eigenfunctions of a harmonic oscillator, and their perturbed counterparts corresponding to an anharmonically altered oscillator.
In essence, the perturbation transforms the state, to this macroscopic superposition:
where, as shown in Fig. 3(c), and are the corresponding energy eigenmodes of the anharmonic potential. Of note, implies entangling interactions between its subsystems, which are photons. One can imagine this final state as being analogous to: in an equivalent spinor representation, where the spin rotation by the angle, models turning on .
In practice, a probabilistically triggered, time-dependent analogue of the above perturbation should also realize the same effect; however, this perturbation should follow the following hierarchy of timescales, such that the modes, and adiabatically evolve into and respectively, and no transitions are induced:
where is the energy difference between the modes, describes the perturbation strength, and is the timescale for carrying out photon number or energy measurements for detecting the inter-particle entanglement of the final state.
Any realization of this Gedankenexperiment will have two requisites, namely: (A) an “ancillary” quantum mechanical subsystem has to be coupled to the entangling system; and (B) a quantum measurement has to be made on this subsystem—as are required for entangling photons that have never interacted in the past [88]. Hence, I have argued that mode-entangled states can be transformed into particle-entangled states by anharmonically and controllably—by a quantum mechanical random process—deforming an initially harmonic potential.
Conclusions.—To summarize, I have examined the equivalence of entanglement between optical modes and entanglement between optical photons that are described by such modes, and have concluded that it is, indeed, possible to transform the former into the latter; usually, the latter is harder to realize and is more useful than the former. Specifically, I have shown that systems having four or more spatial modes, and two or more input unentangled photons—such as non-Abelian, quantum holonomic systems [89, 90, 91] in which light field modes are in photon number entangled relationships with each other [74]—can be used to nonlocally entangle photons. Highly entangled optical systems can be used to construct modular, programmable platforms that are more readily interfaced and integrated with existent long-distance quantum communication networks, which already employ optical photons encoded as flying qubits. Additionally, a future avenue of fundamental research might entail placing the results of this Letter on a firmer mathematical grounding, and extending them to quantum systems found in other branches of physics.
Note added.—Shortly before submitting this paper, I became aware of a recently posted arXiv preprint [92], by the Zeilinger group in Vienna, that sheds new light on the subjective nature of path information, specifically, the “which-way information” in three-particle interferometry. This path information plays a key role in how spatial, “mode-entangled,” or path-entangled superpositions of photons are described.
Acknowledgements.
Acknowledgments.—I would like to thank Chandra Raman for illuminating discussions on the differences between entangled modes and entangled particles, as well as for kindly providing key references on this topic.Author Contributions.—A.B. conceptualized this work, carried out all the calculations and derivations, prepared all the figures, and wrote the manuscript.
Data Availability.—This publication reports a theoretical work, and consequently, no experimental research data were created or analyzed, during the course of this research. In particular, the author confirms that all the relevant information—namely, theoretical calculations, derivations, and accompanying results—that support the findings of this work are available within the article, and its end matter and supplementary material.
Appendix A End Matter
Entanglement measurement contexts in real and Hilbert spaces.—To distinguish between the contextual natures of photon-photon and mode-mode entanglement, consider a quantum optical re-interpretation of the E.P.R.B. Gedankenexperiment, as shown in Fig. 3(a). If we regard the horizontally and vertically polarized, single-photon states, and as the basis states—for example, and —then the corresponding, unitarily transformed states at the measurement station A (orientation angle, ; axis of linear analyzer, ) are:
and similarly, the unitarily transformed states at the measurement station B (orientation angle, ; axis of linear analyzer, ) are:
One can use the above expressions to derive Eq. (2) from Eq. (1), and verify that:
where and are the single-particle, detection probabilities; are the two-particle, joint, detection probabilities; is the coefficient of correlation; and .
Let us now consider two axes of measurement at the measurement station A, and , and two axes of measurement at the measurement station B, and , such that:
and, as can be seen from Fig. 3(a):
Therefore, the well-known algebraic sum of four correlation coefficients, —involving four measurements in four distinct orientations—can be written as:
Figure 3(b) graphically summarizes all these results. Additionally, the entropies of entanglement are found to have no dependence on the relative orientations of the analyzers in space. In contrast, as shown in Fig. 3(c), a rotation in Hilbert space can modulate the expansion coefficients, and, therefore, the degree of entanglement of a mode-mode entangled state. Recently, an approach based on non-Abelian holonomy has been devised to access and tune the angle of this rotation, in a deterministic fashion in a real-world, laboratory setting [74]. Identical results are obtained in Figs. 3(b) and 3(c), if the Rényi entropies are computed, instead of the von Neumann entropies.
Finally, consider the measurement context, as realized by the two-atom, four-momentum-mode interferometer [77] shown in the inset of Fig. 3(d), for detecting the entanglement between the modes of neutral atoms. The four modes are made to interfere two-by-two at two distinct spatial locations, such that the joint detection probabilities are directly accessible and measurable. Specifically, the deflection and mixing are achieved by Bragg diffraction.
Effectively, this interferometer transforms the input state [77]:
where and are the two atomic linear momenta, and and label the two atoms moving in opposite directions, to the output state [77]:
where and are the phase differences between the laser beams forming the two Bragg splitters, and , , , and are the four output modes. Notice that this output state describes genuine inter-modal entanglement, and that all the detection probabilities are directly deducible from the expansion coefficients.
An analysis similar to the one before gives the following expression for the correlation coefficient:
where are the two-atom, joint, detection probabilities. The entropies of entanglement of and are independent of the relative interferometric output phase, , as such phases are relative phases in real space. These results are shown in Fig. 3(d). In essence, the problems of mode basis transformation—pertaining to quantifying the degree of entanglement—and of mutually commuting components of linear momentum along multiple axes—pertaining to quantifying the Bell nonlocality—are solved by an interferometric arrangement with tunable relative phases.